OpenStudy (animusleviosaa):

In a sample of 50 households, the mean number of hours spent on social networking sites during the month of January was 45 hours. In a much larger study, the standard deviation was determined to be 8 hours. Assume the population standard deviation is the same. What is the 95% confidence interval for the mean hours devoted to social networking in January?

OpenStudy (animusleviosaa):

The 95% confidence interval ranges from 8 to 45 hours. <--- not this one. The 95% confidence interval ranges from 40.13 to 45.78 hours. The 95% confidence interval ranges from 43.87 to 46.13 hours. The 95% confidence interval ranges from 42.78 to 47.22

jimthompson5910 (jim_thompson5910):

For 95% confidence interval, the critical z value is z = 1.96 approx. This is something where it's found in a table or you use a calculator Now you'll use these formulas to get the lower bound (L) and the upper bound (U) of the confidence interval \[\Large L = \bar{x} - z*\frac{s}{\sqrt{n}}\] \[\Large U = \bar{x} + z*\frac{s}{\sqrt{n}}\] In this case, \[\Large \bar{x} = 45\] \[\Large z \approx 1.96\] \[\Large s = 8\] \[\Large n = 50\]

OpenStudy (animusleviosaa):

Okay, and when I solve those formulas what would I do?

jimthompson5910 (jim_thompson5910):

Let's compute the lower bound L \[\Large L = \bar{x} - z*\frac{s}{\sqrt{n}}\] \[\Large L = 45 - 1.96*\frac{8}{\sqrt{50}}\] \[\Large L = ???\] I'll let you do the last step. Tell me what you get.

OpenStudy (animusleviosaa):

43.04 * 8/sqrt50

OpenStudy (animusleviosaa):

then sqrt of 50 is 7.07106781

OpenStudy (animusleviosaa):

so we would have 43.04 * 1.13137085

jimthompson5910 (jim_thompson5910):

You don't subtract yet. Subtraction is the last thing you do

jimthompson5910 (jim_thompson5910):

follow PEMDAS

OpenStudy (animusleviosaa):

Oh okay so we would multiply 8/sqrt of 50 by -1.96? how would I do that

jimthompson5910 (jim_thompson5910):

\[\Large L = \bar{x} - z*\frac{s}{\sqrt{n}}\] \[\Large L \approx 45 - 1.96*\frac{8}{\sqrt{50}}\] \[\Large L \approx 45 - 1.96*\frac{8}{7.07106781}\] \[\Large L \approx 45 - 1.96*1.13137085\] \[\Large L \approx 45 - 2.217486866\] \[\Large L \approx 42.782513134\] \[\Large L \approx 42.78\] Hopefully that makes sense?

OpenStudy (animusleviosaa):

Before i saw this I was about to say if it was 42.782513134198987

jimthompson5910 (jim_thompson5910):

good, you got it

jimthompson5910 (jim_thompson5910):

what is the upper bound U? \[\Large U = \bar{x} + z*\frac{s}{\sqrt{n}}\]

OpenStudy (animusleviosaa):

So would the answer be D since it is the only one with that number?

jimthompson5910 (jim_thompson5910):

yes, as practice tell me what you get for the upper bound U

OpenStudy (animusleviosaa):

okay

OpenStudy (animusleviosaa):

before I continue, is this correct? 45+2.217486866

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (animusleviosaa):

okay so would it be 47.217486866

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

side note: `2.217486866` is the approximate margin of error

OpenStudy (animusleviosaa):

Thank you so much :)

OpenStudy (animusleviosaa):

Oh okay, gotcha. Thanks for all of your help

jimthompson5910 (jim_thompson5910):

lower bound = (estimate) - (margin of error) upper bound = (estimate) + (margin of error)

jimthompson5910 (jim_thompson5910):

no problem